A 5 kg brick of 20cm*10cm*8cm dimensionless lying on the largest base. It is now made to stand with the vertical. Find the work done.
 

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Work done = mgh
where h would be half of the longest distance (as to distane to centre of mass.)
tat  would be => 20/2 = 10 cm or => 0.1 meter.
Hence,
 W =mgh/2 = > 5*10*0.1 = > 5 joule.